# I Introduction

### v1

@@ I Introduction
The scalable growth of large-area epitaxial materials is of key importance for any applications involving semiconductor two-dimensional crystals such as transition metal dichalcogenides (TMDs). Among various growth techniques, chemical vapour deposition (CVD) is frequently reported and typically results in a good flatness and structural quality of the achievable crystals due to the high growth temperatures Okada et al. (2014); Wang et al. (2015); Yan et al. (2015); Chae et al. (2017); Carozo et al. (2017); Orsi Gordo et al. (2018); Zhang et al. (2018, 2019a, 2019b); Uchiyama et al. (2019); Shree et al. (2019); Delhomme et al. (2019). While the molecular beam epitaxy (MBE) of TMDs Shimada et al. (1990); Ohuchi et al. (1990); Xenogiannopoulou et al. (2015); Liu et al. (2015); Zhang et al. (2016); Vishwanath et al. (2016); Roy et al. (2016); Onomitsu et al. (2016); Chen et al. (2017); Dau et al. (2017); Nakano et al. (2017); Fu et al. (2017); Wei et al. (2017); Poh et al. (2017, 2018); Vishwanath et al. (2018); Jeong et al. (2018); Dau et al. (2019); Ogorzałek et al. (2020); Seredyński et al. (2022) in an early stage of development, its relatively low growth temperatures offer potential for application on diverse substrate structures, including those containing fragile semiconductor devices. Recently, large, homogeneous structures with excellent optical properties have been successfully grown using MBE Pacuski et al. (2020); Ludwiczak et al. (2021); Połczyńska et al. (2023); Ludwiczak et al. (2024); Vergnaud et al. (2024) thanks to use of hexagonal boron nitride (h–BN) substrates. Despite these advancements, epitaxial techniques face challenges related to the low mobility of transition metals on the substrate surface Mortelmans et al. (2020); Seredyński et al. (2022), which leads to the formation of small grains that later merge into continuous layersFu et al. (2017); Pacuski et al. (2020). Consequently, epitaxial TMD layers suffer from dislocations resulting from grain mergingZhang et al. (2018); Londoño-Calderon et al. (2022) and high amount of point defects Poh et al. (2018). Both grain boundaries and point defects strongly influence the photoluminescence spectrum, particularly by broadening the excitonic emission lines Vergnaud et al. (2019); Pacuski et al. (2020); Kucharek et al. (2023) and significantly reducing the lifetime of the excitonic states by enabling non–radiative relaxation channels Połczyńska et al. (2023); Oreszczuk et al. (2024).
@@
The narrowest exciton lines of MBE-TMDs are obtained for MoSe2 monolayers grown on mechanically exfoliated h-BN Pacuski et al. (2020). A single wafer used in such growth typically contains several flakes of varying thickness on a Si/SiO2 substrate. Alternatively, using Metal-Organic Chemical Vapor Deposition (MOCVD) it is possible to produce wafer-scale h–BN with homogenous thicknesses which acts as an almost equally efficient substrate for MBE growth of MoSe2 Ludwiczak et al. (2021, 2024). In both cases, the resulting TMD structures exhibit high spatial homogeneity in the shape of the photoluminescence spectra, both within individual h-BN flakes and across macroscopic distances on the scale of centimeters for a given sample Pacuski et al. (2020). Test of this homogeneity is limited by the optical spatial resolution, as photoluminescence can only resolve features (inhomogeneities) larger than the diffraction limit, determined by the excitation beam’s wavelength and the numerical aperture of the setup. Consequently, optical measurements involve averaging over a micrometer-sized spot, while the sizes of the MBE–grown MoSe2 grains typically remain below the diffraction limit, in the range of tens of nanometers. It is important to note that the homogeneity of the shape of the photoluminescence spectra across such MBE-grown MoSe2 sample does not entail similar homogeneity in terms of the photoluminescence intensity.
@@ II Samples
III Optical spectra
@@
IV Light intensity comparison
@@
V Interferences effects resulting from h–BN thickness
@@
VI The PL and Raman emission efficiency

### v2

@@ I Introduction
The scalable growth of large-area epitaxial materials is of key importance for any applications involving semiconductor two-dimensional crystals such as transition metal dichalcogenides (TMDs). Among various growth techniques, chemical vapour deposition (CVD) is frequently reported and typically results in a good flatness and structural quality of the achievable crystals due to the high growth temperatures regarding WS2 Okada et al. (2014); Carozo et al. (2017); Orsi Gordo et al. (2018), MoS2 Wang et al. (2015); Yan et al. (2015); Chae et al. (2017); Zhang et al. (2019a); Uchiyama et al. (2019); Shree et al. (2019) or WSe2 Zhang et al. (2018, 2019b); Delhomme et al. (2019). Although molecular beam epitaxy (MBE) of TMDs (WSe2 Jeong et al. (2018); Liu et al. (2015); Nakano et al. (2017), MoS2 Fu et al. (2017), MoTe2 Vishwanath et al. (2016); Roy et al. (2016); Vishwanath et al. (2018); Ogorzałek et al. (2020); Seredyński et al. (2022) or MoSe2 Shimada et al. (1990); Ohuchi et al. (1990); Xenogiannopoulou et al. (2015); Liu et al. (2015); Zhang et al. (2016); Vishwanath et al. (2016); Roy et al. (2016); Onomitsu et al. (2016); Chen et al. (2017); Dau et al. (2017); Wei et al. (2017); Poh et al. (2017, 2018); Jeong et al. (2018); Dau et al. (2019)) is still in an early stage of development, its relatively low growth temperatures offer potential for application on diverse substrate structures, including those containing fragile semiconductor devices and ultra-flat surfaces with low sticking coefficients. Recently, large, homogeneous structures with excellent optical properties have been successfully grown using MBE Pacuski et al. (2020); Ludwiczak et al. (2021, 2024); Vergnaud et al. (2024) thanks to use of hexagonal boron nitride (h–BN) substrates. Despite these advancements, epitaxial techniques face challenges related to the low mobility of transition metals on the substrate surface Mortelmans et al. (2020); Seredyński et al. (2022), which first leads to the formation of small grains that later merge into continuous layers Fu et al. (2017); Pacuski et al. (2020). Consequently, epitaxial TMD layers suffer from dislocations resulting from grain mergingZhang et al. (2018); Londoño-Calderon et al. (2022) and high amount of point defects Poh et al. (2018). Both grain boundaries and point defects strongly influence the photoluminescence spectrum, particularly by broadening the excitonic emission lines Vergnaud et al. (2019); Pacuski et al. (2020); Kucharek et al. (2023) and significantly reducing the lifetime of the excitonic states by enabling non–radiative relaxation channels Połczyńska et al. (2023); Oreszczuk et al. (2024).
@@
The narrowest exciton lines of MBE-TMDs are obtained for MoSe2 monolayers grown on mechanically exfoliated h-BN Pacuski et al. (2020); Vergnaud et al. (2024). A single wafer used in such growth typically contains several flakes of varying thickness on a Si/SiO2 substrate. Alternatively, using Metal-Organic Chemical Vapor Deposition (MOCVD) it is possible to produce wafer-scale h–BN with homogenous thicknesses which acts as an almost equally efficient substrate for MBE growth of MoSe2 Ludwiczak et al. (2021, 2024). In both cases, the resulting TMD structures exhibit high spatial homogeneity in the shape of the photoluminescence spectra, both within individual h-BN flakes and across macroscopic distances on the scale of centimeters for a given sample Pacuski et al. (2020). Test of this homogeneity is limited by the optical spatial resolution, as photoluminescence can only resolve features (inhomogeneities) larger than the diffraction limit, determined by the excitation beam’s wavelength and the numerical aperture of the setup. Other techniques, such as cathodoluminescence Zheng et al. (2017) performed on monolayers or scanning near–field optical microscopy Kim and Kim (2021), come as sample-invasive or very challenging experimentally. Consequently, most of the optical measurements involve signal–averaging over a micrometer-sized spot, while the sizes of the MBE–grown MoSe2 grains typically remain below the diffraction limit, in the range of tens of nanometers.
@@ II Samples
III Experimental techniques
IV Optical spectra
@@
V Light intensity comparison
@@
VI Interferences effects resulting from h–BN thickness
@@
VII The PL and Raman emission efficiency